a certain electromagnet consist of a solenoid 5.0cm long with 200 turns of wire wound on a soft iron core that intensifies the field 130 times find b when the current is 0.3a
1
Expert's answer
2021-07-20T14:54:58-0400

l=0.05m

N=200

"I=0.3A\\\mu_r=130\\\mu_0=4\pi\times10^{-7}Wb/Am"

"B=\mu _0\mu_rni"

"n=\frac{N}{l}=\frac{200}{0.05}=4000"


"B=4\pi\times10^{-7}\times130\times4000\times0.3=0.20T"